Vernon Electronics is considering investing in manufacturing equipment expected to cost $370,000. The equipment has an estimated useful life of four years and a salvage value of $ 21,000. It is expected to produce incremental cash revenues of $185,000 per year. Vernon has an effective income tax rate of 40 percent and a desired rate of return of 14 percent. (PV of $1 and PVA of $1) (Use appropriate factor(s) from the tables provided.) Required Determine the net present value and the present value index of the investment, assuming that Vernon uses straight-line depreciation for financial and income tax reporting. Determine the net present value and the present value index of the investment, assuming that Vernon uses double-declining-balance depreciation for financial and income tax reporting. Determine the payback period and unadjusted rate of return (use average investment), assuming that Vernon uses straight-line depreciation. Determine the payback period and unadjusted rate of return (use average investment), assuming that Vernon uses double-declining-balance depreciation. (Note: Use average annual cash flow when computing the payback period and average annual income when determining the unadjusted rate of return.)
Answer #1

Answer :

Determine the net present value and the present value index of the investment, assuming that Vernon uses straight - line depreciation for financial and income tax reporting.

Cost 370000
Salvage value 21000
Life 4 years
Depreciation = (370000 - 21000)/4 87250
Incremental cash revenues 185000
Depreciation -87250
97750
Less: Tax 40% 39100
EAT 58650
Add : Depreciation 87250
Cash Flow 145900

NPV = - 370000+145900*(PVIFA 14%, 4 years)+21000*(PVIF 14%,4 yrs)

NPV= -370000+145900*(2.912)+21000*(0.592)

NPV = 67292.80

Present value index = (67292.80 + 370000)/370000 = 1.18

Determine the net present value and the present value index of the investment, assuming that Vernon uses double -declining - balance depreciation for financial and income tax reporting.

DDB Schedule

2* Straight - line depreciation rate * Book value at the beginning of the year

2*1/4* Book value

Year Beginning Book Value Depreciation Ending Book Value
1 370000 185000 185000
2 185000 92500 92500
3 92500 46250 46250
4 46250 23125 23125
Year Cash Rev. Depre Net Rev. Tax @40% Depre + EAT Cash flow PVF @14% PV
0 -370000 1 -370000
1 185000 185000 0 0 185000 185000 0.877 162245
2 185000 92500 92500 37000 148000 148000 0.769 113812
3 185000 46250 138750 55500 129500 129500 0.674 87283
4 185000 23125 161875 64750 120250 120250 0.592 71188
NPV 64528

Present value index = (64528 + 370000)/370000 = 1.17

Determine the payback period and unadjusted rate of return (using average investment), assuming that Vernon uses straight - line depreciation.

Year Cash Flow Cumm. Cash flow
0 -370000 -370000
1 145900 -224100
2 145900 -78200
3 145900
4 145900

Payback period = 2yrs + 78200/145900 =2.54 years

Unadjusted rate of return = 58650/(0+370000)/2 = 31.7%

Determine the payback period and unadjusted rate of return (use average investment), assuming that Vernon uses double - declining - balance depreciation

Average income = (0+55500 + 83250 + 97125 ) / 4 = 58968.75

Average investment = 370000 / 2 = 185000

Unadjusted rate of return =58968.75 / (0 + 370000 )/2 = 31.88%

Average cash flow = (185000 +148000 +129500 + 120250 )/4 = 145687.50

Payback period = 370000/145687.50 = 2.54 years.